OpenStudy (anonymous):

Here are a few questions, I would appreciate anything anyone can help me with, I'm required to show work or I won't get points. I can try to return a favor if possible. 6.) 3(2x+4)=5(x+7) 7.) 9|8-4x|=54 8.) 15x+57(less than or equal to sign) 129 9.) -10(less than or equal to sign) 2x+6<12 10.) 2|x-5|>3 11.) -2|4x-5|(less than or equal to sign) 12 THANK YOU SO MUCH! (:

OpenStudy (anonymous):

@dan815 @hba

OpenStudy (anonymous):

@Brainybeauty

OpenStudy (mathmale):

Apply the distributive property of multiplication to both sides of your equation, 3(2x+4)=5(x+7). To get you started off: 3(2x+4) = 7x + 12.

OpenStudy (anonymous):

1) x= 23

OpenStudy (anonymous):

mathmale meant 6x+12 not 7x+23

OpenStudy (mathmale):

Thank you, @bjan. Of course you were right; I typed 7x+12 where I meant 6x+12. 3(2x+4)=5(x+7) is an equation that at first glance appears to be solvable for x. 6x+12 = ??? Please finish the right side. Then consolidate the x terms on the left and the constants on the right, and solve for x.

OpenStudy (anonymous):

I'm not trying to be rude or rush you, but I posted this hours ago because I have to have this submitted by midnight so if you can't just give me the work and answers then I don't have time. but, thank you. I appreciate your willingness to help. Please don't take this comment in a rude way! (:

OpenStudy (mathmale):

Would like to help you, but only in such a way that you learn by doing how to solve problems yourself. Let me know if you change your mind.

OpenStudy (anonymous):

I understand, I just don't have the time. It's 6 minutes til midnight.